DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, October 16, 2002 Padelford C-36 3:45 pm

Singularity of mean curvature flow of symplectic surfaces in a Kahler-Einstein surface

Jingyi Chen (University of British Columbia)

A compact symplectic surface (dim_{ℝ} = 2) remains symplectic along mean curvature flow in a Kahler-Einstein surface (dim_{ℂ} = 2). In general the flow develops singularity in finite time. According to the blow-up rate of the second fundamental form, a singular point is classified into type I or type II singularity. We will discuss the tangent cones at these singular points.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.